OpenStudy (anonymous):

Which percent is equivalent to 5 1/20 ? A. 5.05% B. 5.12% C. 505% D. 520%

OpenStudy (prettygirl_shynice):

ill explain

OpenStudy (anonymous):

k please do

OpenStudy (prettygirl_shynice):

okay

OpenStudy (prettygirl_shynice):

5 1/20=5.20+1=101

OpenStudy (anonymous):

aso

OpenStudy (tkhunny):

1 = 100% 5 = 500%

OpenStudy (prettygirl_shynice):

okay

OpenStudy (tkhunny):

1/2 = 50% 1/4 = 25% 1/5 = 20% 1/10 = 10% 1/20 = 5% Are we getting anywhere?